OpenStudy (anonymous):

Gary and Michelle are taking lunch orders for their office training session. A turkey club sandwich costs $3.75 and a roast beef sandwich costs $3.50. They collected $36.25 from the group of 10 people. How many people ordered Turkey???

OpenStudy (anonymous):

Here you go: you make 2 equations in these types of problems Lets have T = # of turkey sandwiches and R = # of roast beef sandwiches 3.75T + 3.50R = 36.25 second equation is T + R = 10, which means T = 10-R Then you substitute the value of T into the first equation...so 3.75 (10 - R) + 3.50R = 36.25, and then you solve and get R = 5, which means T also = 5, because 10-5 is 5 :)

OpenStudy (anonymous):

Thank you so much!!!! @Kim21

OpenStudy (anonymous):

your welcome :)